# Dynamic survival analysis: modelling the hazard function via ordinary differential equations

### v3

@@ Abstract
The hazard function represents one of the main quantities of interest in the analysis of survival data. We propose a general approach for parametrically modelling the dynamics of the hazard function using systems of autonomous ordinary differential equations (ODEs). This modelling approach can be used to provide qualitative and quantitative analyses of the evolution of the hazard function over time. Our proposal capitalises on the extensive literature of ODEs which, in particular, allow for establishing basic rules or laws on the dynamics of the hazard function via the use of autonomous ODEs. We show how to implement the proposed modelling framework in cases where there is an analytic solution to the system of ODEs or where an ODE solver is required to obtain a numerical solution. We focus on the use of a Bayesian modelling approach, but the proposed methodology can also be coupled with maximum likelihood estimation. A simulation study is presented to illustrate the performance of these models and the interplay of sample size and censoring. Two case studies using real data are presented to illustrate the use of the proposed approach and to highlight the interpretability of the corresponding models. We conclude with a discussion on potential extensions of our work and strategies to include covariates into our framework.
@@ 1 Introduction
Survival analysis represents a classical area of Statistics which is concerned with the analysis of times to event, potentially subject to censoring. Survival analysis methods have been applied in a number of areas, including medicine, epidemiology, genetics, engineering, and biology, to name but a few. The survival function and the hazard function represent two quantities of interest in this area. The survival function provides information about the probability that an individual or population will survive beyond a certain time point. On the other hand, the hazard function represents the instantaneous failure rate at each time point. From a more practical perspective, the hazard function can be interpreted as a quantification of the risk of the event of interest occurring at a specific time, and the evolution of such risk over time.
@@
Several types of estimators of the hazard function have also been proposed. In the frequentist framework, methods based on kernel density estimation and other data-driven smoothing methods have been proposed to estimate the hazard function in the presence of censoring (see Rebora et al., 2014 for a review on this literature). From a Bayesian nonparametric perspective, several types of estimators of the hazard function based on gamma processes (Kalbfleisch, 1978), infinite mixtures (Kottas, 2006), piecewise processes (Arjas and Gasbarra, 1994) have been studied (see Ibrahim et al., 2004 for an overview of Bayesian survival methods in survival analysis). Although flexible, all of these methods lack interpretability, as the shape of the hazard function depends on the number of knots, number of components in the mixture, or smoothing parameters. Indeed, in many cases the estimated hazard may exhibit a wiggly behaviour (McKeague and Tighiouart, 2000; Hess et al., 1999; Kottas, 2006), which complicates the interpretation of such estimates. Another type of estimators of the hazard function, which enjoys some popularity in practice, are piece-wise constant estimators (Friedman, 1982). These estimators impose the condition of constant hazard rates within intervals, and their flexibility depends on the number of intervals and their location. Parametric approaches are often criticised as the shape of the estimated hazard function is restricted to a few possibilities, depending on the model. For example, the hazard function associated with the Weibull distribution can only be increasing, decreasing or flat, while the hazard function associated with the lognormal distribution can only be unimodal (up-then-down). Moreover, the value of the Weibull hazard function at time $t=0$ is either 00 or $\infty$, which is an unrealistic assumption. There exist three-parameter flexible distributions with positive support, such as the power generalised Weibull and generalised gamma, which can capture the basic hazard shapes (increasing, decreasing, unimodal, and bathtub). However, these distributions also impose the condition that the hazard function at time $t=0$ is restricted to be either 00 or $\infty$ (see Rubio et al., 2019 for a review on distributions with flexible hazard functions). Modelling the log-hazard (or log-cumulative-hazard) function using splines has also been considered as an alternative method. However, this requires using a large number of parameters to be able to capture complex shapes, and a careful penalised estimation to obtain smooth curves (see Eletti et al., 2022 for an overview on these methods). More recently, Tang et al. (2023) proposed a framework for modelling the dynamic change of the cumulative hazard function through an ordinary differential equation, but focused on the inclusion of covariates through this formulation based on known hazard structures. In a more particular setting, Tang et al. (2022) consider a general ODE form for the cumulative hazard function (allowing for the inclusion of covariates), and estimate it through a feed-forward neural network with the cumulative hazard as an input. Related approaches have been recently explored in Danks and Yau (2022), who used neural network-based ODEs for modelling cumulative distribution functions. Indeed, these methods tend to focus on adding flexibility to the model, rather than interpreting the hazard or cumulative hazard functions.
@@
We propose a novel approach for modelling the dynamics of the hazard function using systems of first order ODEs, taking the hazard function as a (positive) state variable of the system of ODEs. This modelling approach can also lead to very flexible hazard functions, depending on the system of ODEs employed to characterise the dynamics of the hazard function. This framework is of particular interest to users aiming at understanding and interpreting the evolution of the hazard function over time, in contrast to scenarios where the only aim is to estimate the survival function at specific time points. That is, the proposed approach is relevant in cases where the user is interested in a quantitative and qualitative analysis of the dynamics of the hazard function. We capitalise on the vast literature on ODEs, which provides a number of systems of ODEs with interpretable parameters. In particular, we provide examples using autonomous systems of ODEs which facilitate the interpretation of the dynamics of the hazard function and fill a void in the current literature of hazard-based models. For clarity of exposition and to facilitate a comprehensive analysis of the proposed approach, we will focus on the context without covariates, but we will conclude with a discussion on general strategies for the inclusion of covariates.
@@
In Section 2, we present the model formulation and discuss the advantages of modelling the hazard function through autonomous systems of ODEs. We also compare our formulation against previous models for the cumulative hazard function using ODEs (Tang et al., 2022, 2023). In Section 3, we present two particular models for the hazard function using classical systems of autonomous first order ODEs. Section 4 presents a discussion on the calculation of the likelihood function associated with the proposed modelling approach, covering the cases where the system of ODEs has an analytic solution or where the use of numerical ODE solvers is required to approximate this solution. Section 5 presents a simulation study that illustrates the ability to recover the true values of the parameters and hazard shapes, as well as the effects of sample size and censoring. Section 6 presents two case studies that illustrate the use of the proposed modelling approach using real data. We conclude with a discussion and possible extensions of this work in section 7. Software and real data examples can be found at: https://github.com/FJRubio67/ODESurv for R code, and https://github.com/andreschristen/ODESurv for Python code.
@@ 2 Modelling the hazard function through ODEs
$\begin{cases}{\bf Y}^{\prime}(t)=\psi_{\bm{\theta}}({\bf Y}(t),t),\\ H^{\prime}(t)=h(t),\end{cases}$ (1)
@@
with initial conditions ${\bf Y}(0)={\bf Y}_{0}$ and $H(0)=0$; taking the initial time $t_{0}=0$ as a simplification. Let $D\subset\mathbb{R}^{m+2}$ be a closed rectangle. Assuming that the vector field $\psi_{\bm{\theta}}:D\to\mathbb{R}^{m+1}$ is Lipschitz continuous in ${\bf y}$ for every $t$ ($\psi_{\bm{\theta}}({\bf y},t)$), and that ${\bf Y}_{0}$ is in the interior of $D$, then there exists a unique solution of the above initial value problem on a vicinity of the initial value (for details seePo-Fang and Yasutaka (1999), for example). The vector field $\psi_{\bm{\theta}}$, its domain $D$, and the initial condition ${\bf Y}_{0}$ are such that the solution for $h(t)$ is positive for all $t>0$, and for any parameter value $\bm{\theta}\in\Theta\subset{\mathbb{R}}^{d}$. That is, (1) is a family of systems of ODEs for which one state variable is the hazard function $h$, leading to a family of hazard functions defined by $\bm{\theta}$ and ${\bf Y}_{0}$. The cumulative hazard function $H$ is also included in the formulation (1), and is thus obtained in the solution of the system. The initial condition must satisfy $h(0)\geq 0$, but it is otherwise arbitrary. The initial conditions in (1) could include $h(0)=h_{0}>0$ indicating that the hazard function takes a non-negative, finite, value at $t=0$, and that all individuals are alive at the start of the follow-up. There is no unique rule for determining the number of hidden states in the ODE system (1). In practice, the number of hidden states to be included in (1) may depend on the biological interpretation of the system, or additional states may either help add flexibility or additional effects to the model. In Section 3, we present an example of an ODE system with two states which represent the interaction between the hazard function associated to a disease and the response resulting from an intervention (treatment) on the population. In section 7, we present an example of a system with three states, where one of the states is included based on a biological interpretation (similar to the aforementioned ODE system), and the third state is included to model time-delay effects.
@@
where ${\bf x}\in{\mathbb{R}}^{p}$ are the available covariates. This formulation aims at modelling the dynamics of the cumulative hazard function, which implicitly means modelling the hazard function $h(t;{\bf x})={H^{\prime}}(t;{\bf x})$, rather than obtaining it as a solution, in contrast to our formulation (1). Moreover, our formulation (1) produces the hazard function and the cumulative hazard function as solutions of a system of ODEs, in contrast to (2). Finally, formulation (1) allows for including hidden states variables through the inclusion of additional equations in the system. However, some care is needed to guarantee identifiability of parameters, as is the case of any parametric model. On the other hand, note that by setting $h(t\mid\bm{\theta})=\Psi(H(t\mid\bm{\theta}),t)$ and $\Psi(H(t_{0}\mid\bm{\theta}),t_{0})={\bf Y}_{0}=h_{0}$, we obtain an (theoretical) equivalence between our formulation (1) and formulation (2), restricted to survival models with differentiable hazard functions and models without additional hidden states. Tang et al. (2022) propose to define $\Psi$ through a neural network, with several parameters to estimate. Then, the formulation (2) is not seen by Tang et al. (2022) as an approach to model the (cumulative) hazard function, but simply as a device to use a neural network to construct it.
@@
In some cases, the solution to the system of ODEs (1) is analytic, as we will show in our example in Section 3.1.1. However, this is more the exception than the rule, as many systems of ODEs do not have an analytic solution. In such cases, one can obtain the solution at specific time points $h(t_{i})$ by using an appropriate ODE numerical solver, given the vector field $\psi_{\bm{\theta}}$ (colloquially known as the “right hand side” or rhs), the initial conditions ${\bf Y}(0)={\bf Y}_{0},H(0)=0$ and the time points $t_{1},\ldots,t_{n}$ where the solution must be approximated. Fortunately, the theory of the numerical analysis for systems of ODEs, i.e. initial value problems, is quite robust (seeButcher (2016), for example). Solver implementations include sophisticated implicit step algorithms that dynamically choose methods for stiff and non-stiff systems (e.g. the “Livermore Solver for Ordinary Differential Equations with Automatic method switching”). Moreover, easy-to-use wrappers of these solvers are available both in R (deSolve, Soetaert et al., 2010), and in Python (scipy.integrate.solve_ivp, Jones et al., 2001). The examples presented in Section 3 are implemented both in R and in Python using these solvers.
@@ 2.1 Autonomous systems of ODEs for modelling the hazard function
If the vector field in the system of ODEs (1) does not explicitly depend on $t$, it is said to be an autonomous system. That is,
$\begin{cases}{\bf Y}^{\prime}(t)=\psi_{\bm{\theta}}({\bf Y}(t)),\\ H^{\prime}(t)=h(t).\end{cases}$ (3)
with ${\bf Y}(0)={\bf Y}_{0}$ and $H(0)=0$. The functions $q_{1}(t),\ldots,q_{m}(t)$ are hidden state variables, adding flexibility and richness to the modelling of the hazard function $h$. The simplest case, without hidden state variables, contains only ${\bf Y}=h$. In Section 3.1 we present an example of the latter and in Section 3.2 and example of the former, with one hidden state variable $q$.
Note that any given smooth hazard function $\tilde{h}$ may be trivially formulated as the solution of a non-autonomous ODE, with $\psi_{\bm{\theta}}(\tilde{h},t)=\tilde{h}^{\prime}(t)$. It would be interesting to understand when an how any given hazard function (e.g. for the Weibull or Lognormal distribution, see Section 3.1) can be written in terms of an initial value problem with autonomous ODEs as in (3), establishing the corresponding minimal dimension of ${\bf Y}$, with a simple enough $\psi_{\theta}$. In fact, Kunze and Vrscay (1999) show how a target smooth function may be arbitrarily close to the solution of an autonomous ODE, using an $N$ degree polynomial $\psi(x)$, for $N$ sufficiently large. This result in turn shows that the ODE system (1) can be used to produce very flexible hazard functions. However, the principal concern of this paper is not restricted to producing yet another flexible hazard modelling framework, but to explore the use of autonomous systems of ODEs, as in (3), to model the hazard function $h$ with a clear and useful interpretation in terms of the survival problem at hand.
As opposed to time dependent non-autonomous systems, autonomous systems provide qualitative insights regarding the dynamic evolution of the system. Since the times of Isaac Newton, autonomous ODEs have been used to model systems by stating principles or laws, namely, the vector field $\psi_{\bm{\theta}}$, that remain fixed in time. This powerful tool has been successfully used to model all sorts of phenomena through “basic rules”, ranging from a variety of areas such as physics, chemistry, biology, epidemics, and sociology, to name but a few (Borzi, 2022). Our goal is to harness this powerful tool to model hazard functions, for specific problems, based on basic rules dictated by the structure of the system of ODEs. In fact, general qualitative descriptions of hazard functions are already common in survival analysis and reliability (Simes and Zelen, 1985; Royston and Parmar, 2002; Rubio et al., 2019; Meeker et al., 2022). Here we aim to go one step further, turning qualitative assessments into basic interpretable rules, as a tool for modelling hazard functions, as will be explained in the following examples.

### v4

@@ Abstract
The hazard function represents one of the main quantities of interest in the analysis of survival data. We propose a general approach for parametrically modelling the dynamics of the hazard function using systems of autonomous ordinary differential equations (ODEs). This modelling approach can be used to provide qualitative and quantitative analyses of the evolution of the hazard function over time. Our proposal capitalises on the extensive literature of ODEs which, in particular, allow for establishing basic rules or laws on the dynamics of the hazard function via the use of autonomous ODEs. We show how to implement the proposed modelling framework in cases where there is an analytic solution to the system of ODEs or where an ODE solver is required to obtain a numerical solution. We focus on the use of a Bayesian modelling approach, but the proposed methodology can also be coupled with maximum likelihood estimation. A simulation study is presented to illustrate the performance of these models and the interplay of sample size and censoring. Two case studies using real data are presented to illustrate the use of the proposed approach and to highlight the interpretability of the corresponding models. We conclude with a discussion on potential extensions of our work and strategies to include covariates into our framework. Although we focus on examples on Medical Statistics, the proposed framework is applicable in any context where the interest lies on estimating and interpreting the dynamics hazard function.
@@ 1 Introduction
Survival analysis represents a classical area of Statistics which is concerned with the analysis of times to event, potentially subject to censoring. Survival analysis methods have been applied in a number of areas, including medicine, epidemiology, genetics, engineering, and biology, to name but a few. The survival function and the hazard function represent two quantities of interest in this area. The survival function provides information about the probability that an individual or population will survive beyond a certain time point. On the other hand, the hazard function represents the instantaneous failure rate at each time point. From a more practical perspective, the hazard function can be interpreted as a quantification of the risk of the event of interest occurring at a specific time, and the evolution of such risk over time (Rinne, 2014).
@@
Several types of estimators of the hazard function have also been proposed. In the frequentist framework, methods based on kernel density estimation and other data-driven smoothing methods have been proposed to estimate the hazard function in the presence of censoring (see Rebora et al., 2014 for a review on this literature).
From a Bayesian nonparametric perspective, several types of estimators of the hazard function based on gamma processes (Kalbfleisch, 1978), infinite mixtures (Kottas, 2006), piecewise processes (Arjas and Gasbarra, 1994) have been studied (see Ibrahim et al., 2004 for an overview of Bayesian survival methods in survival analysis). Although flexible, all of these methods lack interpretability, as the shape of the hazard function depends on the number of knots, number of components in the mixture, or smoothing parameters. Indeed, in many cases the estimated hazard may exhibit a wiggly behaviour (McKeague and Tighiouart, 2000; Hess et al., 1999; Kottas, 2006), which complicates the interpretation of such estimates.
Another type of estimators of the hazard function, which enjoys some popularity in practice, are piece-wise constant estimators (Friedman, 1982). These estimators impose the condition of constant hazard rates within intervals, and their flexibility depends on the number of intervals and their location.
Parametric approaches are often criticised as the shape of the estimated hazard function is restricted to a few possibilities, depending on the model. For example, the hazard function associated with the Weibull distribution can only be increasing, decreasing or flat, while the hazard function associated with the lognormal distribution can only be unimodal (up-then-down). Moreover, the value of the Weibull hazard function at time $t=0$ is either 00 or $\infty$, which is an unrealistic assumption. There exist three-parameter flexible distributions with positive support, such as the power generalised Weibull and generalised gamma, which can capture the basic hazard shapes (increasing, decreasing, unimodal, and bathtub). However, these distributions also impose the condition that the hazard function at time $t=0$ is restricted to be either 00 or $\infty$ (see Rubio et al., 2019 for a review on distributions with flexible hazard functions).
@@
Modelling the log-hazard (or log-cumulative-hazard) function using splines has also been considered as an alternative method. However, this requires using a large number of parameters to be able to capture complex shapes, and a careful penalised estimation to obtain smooth curves (see Eletti et al., 2022 for an overview on these methods). More recently, Tang et al. (2023) proposed a framework for modelling the dynamic change of the cumulative hazard function through an ordinary differential equation, but focused on the inclusion of covariates through this formulation based on known hazard structures. In a more particular setting, Tang et al. (2022) consider a general ODE form for the cumulative hazard function (allowing for the inclusion of covariates), and estimate it through a feed-forward neural network with the cumulative hazard as an input. Related approaches have been recently explored in Danks and Yau (2022), who used neural network-based ODEs for modelling cumulative distribution functions. Indeed, these methods tend to focus on adding flexibility to the model, rather than interpreting the hazard or cumulative hazard functions.
@@
We propose a novel approach for modelling the dynamics of the hazard function using systems of first order ODEs, taking the hazard function as a (positive) state variable of the system of ODEs. In our context, the word dynamics refers to the description of the evolution of the hazard function over time through systems of ODEs. This modelling approach can also lead to very flexible hazard functions, depending on the system of ODEs employed to characterise the dynamics of the hazard function. For instance, it is well known that autonomous scalar ODEs have monotonic solutions (Wallach, 1948, see also Section 3). Consequently, using a single ODE to model the hazard function can only capture increasing or decreasing shapes of the hazard function. However, by incorporating additional states, it is possible to obtain non-monotonic hazard shapes, as shown in Section 3. These internal states allow us to account for complex interactions, delays, or memory effects (Farkas, 1984) that would not be possible to model with a scalar ODE (see Section 7 for an example). Therefore, the choice of the system of ODEs for modelling the hazard function allows for adding flexibility in an interpretable manner, as the additional hidden states not only enrich the family of hazard function models but they also have a clear dynamic influence on the hazard function itself. The proposed framework is of particular interest to users aiming at understanding and interpreting the evolution of the hazard function over time, in contrast to scenarios where the only aim is to estimate the survival function at specific time points. That is, the proposed approach is relevant in cases where the user is interested in a quantitative and qualitative analysis of the dynamics of the hazard function. We capitalise on the vast literature on ODEs, which provides a number of systems of ODEs with interpretable parameters. In particular, we provide examples using autonomous systems of ODEs which facilitate the interpretation of the dynamics of the hazard function and fill a void in the current literature of hazard-based models. For clarity of exposition and to facilitate a comprehensive analysis of the proposed approach, we will focus on the context without covariates, but we conclude with a discussion on general strategies for the inclusion of covariates.
@@
In Section 2, we present the model formulation and discuss the advantages of modelling the hazard function through autonomous systems of ODEs. We present a discussion on modern tools to verify identifiability of hazard models obtained as a solution of systems of ODEs. We also compare our formulation against previous models for the cumulative hazard function using ODEs (Tang et al., 2022, 2023). In Section 3, we present two particular models for the hazard function using classical systems of autonomous first order ODEs. We provide an interpretation of the additional hidden states and the added flexibility obtained through their inclusion. Section 4 presents a discussion on the calculation of the likelihood function associated with the proposed modelling approach, covering the cases where the system of ODEs has an analytic solution or where the use of numerical ODE solvers is required to approximate this solution. Section 5 presents a simulation study that illustrates the ability to recover the true values of the parameters and hazard shapes, as well as the effects of sample size and censoring. Section 6 presents two case studies that illustrate the use of the proposed modelling approach using real data. We conclude with a discussion and possible extensions of this work in Section 7. Software and real data examples can be found at: https://github.com/FJRubio67/ODESurv for R code, and https://github.com/andreschristen/ODESurv for Python code.
@@ 2 Modelling the hazard function through ODEs
$\begin{cases}{\bf Y}^{\prime}(t)=\psi_{\bm{\theta}}({\bf Y}(t),t),\\ H^{\prime}(t)=Y_{1}(t),\end{cases}$ (1)
@@
with initial conditions ${\bf Y}(0)={\bf Y}_{0}$ and $H(0)=0$; taking the initial time $t_{0}=0$ as a simplification, and $Y_{1}(t)=h(t)$ by definition. Let $D\subset\mathbb{R}^{m+2}$ be a closed rectangle. Assuming that the vector field $\psi_{\bm{\theta}}:D\to\mathbb{R}^{m+1}$ is Lipschitz continuous in ${\bf y}$ for every $t$ ($\psi_{\bm{\theta}}({\bf y},t)$), and that ${\bf Y}_{0}$ is in the interior of $D$, then there exists a unique solution of the above initial value problem on a vicinity of the initial value (for details see Po-Fang and Yasutaka (1999), for example). The vector field $\psi_{\bm{\theta}}$, its domain $D$, and the initial condition ${\bf Y}_{0}$ are such that the solution for $h(t)$ is positive for all $t>0$, and for any parameter value $\bm{\theta}\in\Theta\subset{\mathbb{R}}^{d}$. That is, (1) is a family of systems of ODEs for which one state variable is the hazard function $h$, leading to a family of hazard functions defined by $\bm{\theta}$ and ${\bf Y}_{0}$. The cumulative hazard function $H$ is also included in the formulation (1), and is thus obtained in the solution of the system. The initial condition must satisfy $h(0)\geq 0$, but it is otherwise arbitrary. The initial conditions in (1) could include $h(0)=h_{0}>0$ indicating that the hazard function takes a non-negative, finite, value at $t=0$, and that all individuals are alive at the start of the follow-up.
While there is no general characterisation of ODE systems guaranteeing positive solutions (specifically, for $Y_{1}(t)$), several strategies can ensure non-negativity. One approach involves directly verifying non-negativity through a qualitative analysis of solutions based on the specific structure of the ODE system. This technique is used, for instance, to prove non-negativity in competitive species models, like the Lotka-Volterra equations (Boyce et al., 2021), which we study in Section 3. Indeed, a vast literature exists on population dynamics models using ODEs where solutions are naturally positive (Hirsch et al., 2012).
There are also particular theoretical characterisations of systems of ODEs with positive solutions, which are often more restrictive. For example, Bernstein and Bhat (1999) demonstrates that for autonomous systems (see Section 2.1), if the vector field is essentially non-negative (i.e. entry-wise) and the initial conditions are non-negative, then the solutions are non-negative. Another strategy involves modelling the dynamics of a transformed hazard function instead of directly modelling the hazard function, such as $\tilde{h}(t)=\log h(t)$, which allows for using vector fields that can be negative. This approach is also used to improve computational efficiency and stability of ODE solvers, as discussed in Sections 3 and 6.
Finally, another option consists of “forcing” the solution to become positive (Blanes et al., 2022). This can be done theoretically or numerically by adding a parameter-dependent offset that shifts the solution to the positive quadrant or by discarding parameter values that lead to negative solutions altogether.
In any case, establishing the positivity of state variables in a system of ODEs, as described in (1), needs to be done on a case-by-case basis, considering the characteristics of $\psi_{\bm{\theta}}$ and the initial conditions ${\bf Y}_{0}$.
Regarding the number of state variables $q_{j}$’s, or hidden states, there is no unique rule for determining the number of $q_{j}$’s in the ODE system (1). In practice, the number of hidden states to include in (1) may depend on the biological interpretation of the system. Additional states can also be added to enhance the model’s flexibility or incorporate additional effects. In Section 3, we present an example of an ODE system with two states which represent the interaction between the hazard function associated to a disease and the response resulting from an intervention (treatment) on the population. In Section 7, we present an example of a system with three states, where the third state is included to model time-delay effects.
The increased flexibility obtained through additional hidden states is appealing, but as with any parametric model, ensuring parameter identifiability is crucial for making inference about the parameters. Establishing conditions to guarantee identifiability of the solution of systems of ODEs has received a fair amount of attention in the last couple of decades. Studies like Miao et al. (2011) and Qiu et al. (2022) provide conditions to verify identifiability of parameters for families of linear and non-linear systems of ODEs. Additionally, the Julia package ‘StructuralIdentifiability.jl’ implements state-of-the-art numerical methods to detect non-identifiability. Therefore, we encourage users of our framework to take advantage of these tools to verify identifiability of the parameters of any new system of ODEs. Nonetheless, the literature on systems of ODEs is vast, and the identifiability of many of the classical models (such as those presented in Section 3) has already been established.
@@
where ${\bf x}\in{\mathbb{R}}^{p}$ are the available covariates. This formulation aims at modelling the dynamics of the cumulative hazard function, which implicitly means modelling the hazard function $h(t;{\bf x})={H^{\prime}}(t;{\bf x})$, rather than obtaining it as a solution, in contrast to our formulation (1). Moreover, our formulation (1) produces the hazard function and the cumulative hazard function as solutions of a system of ODEs, in contrast to (2). Finally, formulation (1) allows for including hidden states variables through the inclusion of additional equations in the system.
On the other hand, note that by setting $h(t\mid\bm{\theta})=\Psi(H(t\mid\bm{\theta}),t)$ and $\Psi(H(t_{0}\mid\bm{\theta}),t_{0})={\bf Y}_{0}=h_{0}$, we obtain an (theoretical) equivalence between our formulation (1) and formulation (2), restricted to survival models with differentiable hazard functions and models without additional hidden states. More importantly, Tang et al. (2022) propose to define $\Psi$ through a neural network, with several parameters to estimate. Then, the formulation (2) is not seen by Tang et al. (2022) as an approach to model the (cumulative) hazard function, but simply as a device to use a neural network to construct it.
@@
In some cases, the solution to the system of ODEs (1) is analytic, as we will show in our example in Section 3.2. However, this is more the exception than the rule, as many systems of ODEs do not have an analytic solution. In such cases, one can obtain the solution at specific time points $h(t_{i})$ by using an appropriate ODE numerical solver, given the vector field $\psi_{\bm{\theta}}$ (colloquially known as the “right hand side” or rhs), the initial conditions ${\bf Y}(0)={\bf Y}_{0},H(0)=0$ and the time points $t_{1},\ldots,t_{n}$ where the solution must be approximated. Fortunately, the theory of the numerical analysis for systems of ODEs, i.e. initial value problems, is quite robust (see Butcher (2016), for example). Solver implementations include sophisticated implicit step algorithms that dynamically choose methods for stiff and non-stiff systems (e.g. the “Livermore Solver for Ordinary Differential Equations with Automatic method switching”). Moreover, easy-to-use wrappers of these solvers are available both in R (deSolve, Soetaert et al., 2010), and in Python (scipy.integrate.solve_ivp, Jones et al., 2001). The examples presented in Section 3 are implemented both in R and in Python using these solvers.
@@ 2.1 Autonomous systems of ODEs for modelling the hazard function
If the vector field in the system of ODEs (1) does not explicitly depend on $t$, it is said to be an autonomous system (Boyce et al., 2021). That is,
$\begin{cases}{\bf Y}^{\prime}(t)=\psi_{\bm{\theta}}({\bf Y}(t)),\\ H^{\prime}(t)=Y_{1}(t).\end{cases}$ (3)
with ${\bf Y}(0)={\bf Y}_{0}$ and $H(0)=0$, and $Y_{1}(t)=h(t)$. The functions $q_{1}(t),\ldots,q_{m}(t)$ are hidden state variables, adding flexibility and richness to the modelling of the hazard function $h$, as shown in the examples in Sections 3 and 7. In fact, Kunze and Vrscay (1999) show how a target smooth function may be arbitrarily close to the solution of an autonomous ODE, using an $N$ degree polynomial $\psi(h)$, for $N$ sufficiently large. This result in turn shows that the ODE system (3) can be used to produce very flexible hazard functions. However, the primary focus of this paper extends beyond simply developing another flexible hazard modeling framework. Instead, it aims to explore the application of autonomous systems of ODEs, as in (3), to model the hazard function $h$ in a manner that offers a clear and useful interpretation concerning the survival problem under consideration.
As opposed to time dependent non-autonomous ODEs, autonomous ODEs provide qualitative insights regarding the dynamic evolution of the system. In Section 3, we present an example illustrating a qualitative analysis of a system of ODEs, which shows how the solutions evolve over time by examining their equilibrium points. Since the times of Isaac Newton, autonomous ODEs have been used to model systems by stating principles or laws, namely, the vector field $\psi_{\bm{\theta}}$, that remain fixed in time. This powerful tool has been successfully used to model all sorts of phenomena through “basic rules”, ranging from a variety of areas such as physics, chemistry, biology, epidemics, and sociology, to name but a few (Borzi, 2022). Our goal is to harness this powerful tool to model hazard functions, for specific problems, based on basic rules dictated by the structure of the system of ODEs. In fact, general qualitative descriptions of hazard functions are already common in survival analysis and reliability (Simes and Zelen, 1985; Royston and Parmar, 2002; Rubio et al., 2019; Meeker et al., 2022). Here we aim to go one step further, turning qualitative assessments into basic interpretable rules, as a tool for modelling hazard functions, as will be explained in the examples in Section 3.